Ronald Roth started his new job as controller with Aerosystems today. Carole, the employee benefits clerk, gave Ronald a packet that contains information on the company’s health insurance options. Aerosystems offers its employees the choice between a private insurance company plan (Blue Cross/Blue Shield), an HMO, and a PPO. Ronald needs to review the packet and make a decision on which health care program fits his needs. The following is an overview of that information. a) The monthly premium cost to Ronald for the Blue Cross/Blue Shield plan will be $53.61. For all doctor office visits, prescriptions, and major medical charges, Ronald will be responsible for 15 percent and the insurance company will cover 85 percent of covered charges. The annual deductible is $720. b) The HMO is provided to employees free of charge. The copayment for doctors’ office visits and major medical charges is $15. Prescription copayments are $10. The HMO pays 100 percent after Ronald’s copayment. There is no annual deductible. c) The POS requires that the employee pay $35.81 per month to supplement the cost of the program with the company’s payment. If Ron uses health care providers within the plan, he pays the copayments as described above for the HMO with no annual deductible. He can also choose to use a healthcare provider out of the network and pay 15 percent of all charges after he pays a $720 deductible. The POS will pay for 85 percent of those covered visits. Ronald decided to review his medical bills from the previous year to see what costs he had incurred and to help him evaluate his choices. He visited his general physician four times during the year at a cost of $180 for each visit. He also spent $76 and $100 on two prescriptions during the year. What annual medical costs will Ronald pay using the sample medical expenses provided if he enrolls in the Blue Cross/Blue Shield plan? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

In the entire text, I could figure out just one question. I am answering that question:

What annual medical costs will Ronald pay using the sample medical expenses provided if he enrolls in the Blue Cross/Blue Shield plan? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Annual medical costs = Annual premiums + Deductible + Coinsurance cost

Annual premium = Monthly premium x 12 = 53.61 x 12 = 643.32

Deductible = 720

Coinsurance cost = Coinsurance percent x Coinsurance expenses

Coinsurance expenses = Number of Physician visits × Cost per visit + Prescription costs = 4 x 180 + (76 + 100) = 896

Hence, Coinsurance cost = 15% x 896 = 134.40

Hence, Annual medical costs = 643.32 + 720 + 134.40 = $ 1,497.72